OpenStudy (anonymous):

show me how to solve this equation please: c^2+(3c)+2=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (mathteacher1729):

This is a quadratic equation. It can be factored into two binomials. In pictures, you want to find values of x where the curve intersects the x axis. (see picture)

OpenStudy (anonymous):

Factorise (split into brackets):\[c^2+3c+2=(c+2)(c+1)=0\]Then, you have two numbers that when multiplied together make 0, so one of them is zero. So, either:\[c+2=0=>c=-2\]or\[c+1=0=>c=-1\]

OpenStudy (mathteacher1729):

To avoid confusion with equal/greater than signs: \Rightarrow \[\Rightarrow\]

OpenStudy (anonymous):

Ah, thanks! I wondered if there was a proper implies sign on her.

OpenStudy (mathteacher1729):

Draw your desired symbol, it gives you the code: http://detexify.kirelabs.org/classify.html Truly, we live in the future!

OpenStudy (anonymous):

Wow!